# Probing Primordial Black Hole Formation from Domain Wall Isocurvature Perturbations: Constraints and Implications

### v2

@@ Abstract
Domain walls are topological defects produced by the spontaneous symmetry-breaking of discrete symmetry during cosmological phase transitions. The horizon-size domain wall can significantly contribute to the energy density in the late-evolution stage. We propose that the density perturbations from the fluctuations in the number density of the horizon-size domain wall could collapse to form primordial black holes. This mechanism becomes effective when the domain wall energy density ratio to that of the radiation reaches about 0.1 in the radiation-dominated Universe. We find that models with $Z_{2}$ symmetry are excluded for interpreting pulsar timing array observations on the nano-Hz gravitational wave background since this model’s domain wall number density fluctuations could lead to an overabundance of the primordial black holes. Moreover, models with $N\sim 10$ domain walls also suffer strong constraints from the overabundance of primordial black holes.
@@ 1 Introduction
It is widely believed that new physics is required to explain various phenomena beyond the Standard Model (SM), including the neutrino mass and mixing [1], dark matter [2], strong CP problem [3], and matter-antimatter asymmetry [4], etc. As the cosmic temperature drops with the expansion of the Universe, the spontaneous breakdown of discrete symmetries during a phase transition in new physics models can generate sheet-like topological defects, the domain walls [5, 6]. The Universe after the phase transition is then randomly divided by them into distinct domains, which represent the clumps of the degenerate vacua. The domain walls tend to stretch in size and the neighboring domains with the same vacua merge to stabilize the network. This evolution can be described by the percolation theory [7], whose fundamental results are based on the existence of a percolation threshold $p_{c}=0.311$ for a simple cubic lattice. The domain wall’s energy density scales as $\rho_{w}\propto t^{-1}$ in the scaling regime, which is much slower than the decay of radiation and dust, and eventually dominates the cosmological evolution. In the domain wall era, the repulsive force from domain walls would lead to an inflation phase of the Universe with the scale factor going as $a\propto t^{2}$. Such an expansion history is not compatible with the observations since it could strongly alter the Big Bang nucleosynthesis (BBN), reduce the time for the galaxy formation, and lead to unacceptably large anisotropies of the cosmic microwave background (CMB) [8]. Nevertheless, the research on the nature of dark energy in the context of domain wall has been continuously conducted by recent works [9, 10, 11].
@@
Discrete symmetries exist widely in new physics models. A $Z_{2}$ symmetry can emerge from the spontaneous breakdown of a gauge group in the grand unified theory (GUT) [12, 13]. The $Z_{2}$ symmetry is also commonly imposed on the two-Higgs-doublet models for the spontaneous CP violation to suppress the Higgs-mediated flavor-changing neutral currents (FCNCs) at tree level [14, 15, 16]. The global $U(1)$ symmetry in the axion models for solving the strong CP problem is explicitly broken to the $Z_{\mathcal{N}}$ symmetry by the QCD instanton effect [17, 15]. Furthermore, to provide a dark matter candidate, a $Z_{2}$ symmetry is often imposed to make the dark matter particle stable [18, 19, 20, 21, 22, 23, 24].
@@
A common solution to the domain wall problem is to introduce a bias potential $V_{\rm bias}$ to break the degenerate vacua [25, 26]. The domain wall network collapses as its surface energy is comparable with the bias potential, $\sigma H\simeq V_{\rm bias}$. The bias potential can be generated by the instanton effects if the global symmetry is anomalous under the framework of the theory, with such examples as the QCD and quantum gravity [15, 25, 26, 27, 28, 29]. It is notable that based on the successful and precise prediction in Ref. [30], Ref. [29] suggest that the recent pulsar timing array (PTA) observations, including the NANOGrav 15-year dataset (NG15) and IPTA-DR2 dataset, strongly support the mechanism of the QCD instanton-induced domain wall annihilations.
@@
Primordial black holes (PBHs) have attracted much attention since it was first suggested by Zel’dovich more than 50 years ago [31]. Contrary to the black holes resulting from the death and collapse of stars in their final evolution stage, PBHs are formed by cosmological processes in the early Universe. In 1971, Hawking proposed the first mechanism of PBH production [32, 33]: the fluctuations of a radiation-dominated Universe can produce some overdense regions. Once the density of the region $\delta=\delta\rho/\bar{\rho}$ exceeds a critical value $\delta_{c}\simeq w$, the overdense region can stop expanding and re-collapse to form PBHs when the region enters the horizon. Cosmic inflation [34, 35, 36] is commonly considered a primary source of density perturbation. The large curvature perturbations generated by the quantum fluctuations of the scalar fields are dominated by the density perturbations once they exit the horizon [36]. For other PBH production mechanisms, we refer to Refs. [37, 38] for a recent review.
@@
In contrast to PBH formation from the collapse of closed domain wall [39, 40, 41, 42, 43, 44], we show in this work that the density perturbation due to the number density fluctuations of horizon-size domain walls in the late stage could lead to the formation of PBHs. The collapse of a closed domain wall needs to satisfy the Schwarzschild criterion, i.e., the closed wall has a Schwarzschild radius near collapse [41]. Our scenario is based on the critical collapse of horizon-size perturbations, as outlined above. We find that if the annihilations of the domain wall are the source of the nano-Hz stochastic gravitational wave background (SGWB) observed recently by PTAs, a significant PBH relic abundance could be produced via the number density fluctuations of the walls before their annihilations.
@@ 2.1 Domain wall equation of state
The energy-momentum tensor of a plane domain wall lying in the $yz$ plane at $x=0$ (neglecting the thickness of the wall) can be estimated as [45, 46]
@@
where $\overline{L}=L/N$ is the average wall separation, in which there exists on the average one domain wall with the surface energy-momentum tensor
@@
The average energy-momentum tensor is determined by $\overline{T}_{\mu\nu}=S_{\mu\nu}/\overline{L}$. Note that since the wall can extend along any direction, we interpret $\overline{L}$ as the correlation length of the wall. The energy density and pressure of the wall are given by temporal and spatial components of the energy-momentum tensor
$\rho_{w}=\overline{T}_{00}=\frac{\gamma^{2}\sigma_{w}}{\overline{L}},\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ \leavevmode\nobreak\ p_{w}=\overline{T}_{ii}=\frac{\sigma_{w}}{3\overline{L}}(\overline{v}_{w}^{2}\gamma^{2}-2).$ (2.5)
@@ 2.2 The evolution of a domain wall
The correlation length $\bar{L}$ (or curvature radius) and the root-mean-square velocity $\bar{v}_{w}$ are two key physical quantities for describing the domain walls. We first adopt the so-called VOS model [47] (see appendix C for more details) to describe the evolution of $\bar{L}$ and $\bar{v}_{w}$ with cosmic expansion, which is given by
@@
where the momentum parameter $k_{w}=0.66\pm 0.04$ and the chopping parameter $c_{w}=0.81\pm 0.04$ are derived from simulations of domain wall evolution in a radiation-dominated Universe [48]. The momentum parameter $k_{w}$ characterizes the curvature acceleration of the wall, while the chopping parameter $c_{w}$ represents the interaction and merger between walls. The damping length is defined as
@@
In the upper panel of Fig. 1, we illustrate the evolution of the domain wall formed at time $t_{s}$ with a symmetry-breaking scale of $v=\sigma_{w}^{1/3}=10^{5}$ GeV. The formation time of the domain wall is determined by $t_{s}=M_{\rm Pl}/(3.32g_{*}^{1/2}T_{s}^{2})$, where $g_{*}(T)\simeq 106.75$ represents the effective degrees of freedom of the plasma, $M_{\rm Pl}=1.2\times 10^{19}$ GeV is the Planck mass, and the domain wall formation temperature is assumed to be $T_{s}=v$. The initial conditions $\bar{L}(t_{s})=0.01t_{s}$ and $\bar{v}_{w}\to 0$ are adopted to solve Eqs. (2.7). We further assume a vacuum pressure $V_{\rm bias}=(10\leavevmode\nobreak\ \rm GeV)^{4}$, with the onset of the bias potential occurring at $t_{\rm on}=0.01t_{\rm ann}$.
The red curves depict the evolution of $\bar{L}(t)$, while the blue curves represent $\bar{v}_{w}$. To emphasize the role of chopping, the parameter $c_{w}$ is set to $0.81$ and $0.61$ for the solid and dashed curves, respectively. We observe that the domain wall rapidly expands to a superhorizon size, $\bar{L}\sim 6t$, driven by curvature acceleration resulting from the wall tension. Subsequently, the correlation length of the wall decreases due to the Hubble friction, chopping effects, and friction from the bias potential. As the cosmic time approaches $t_{\rm ann}$, $\bar{L}(t)$ undergoes a significant reduction, re-entering the horizon before the domain wall annihilates at $t_{\rm ann}$. Initially, the domain walls are generated as a radiation gas, characterized by $\bar{v}_{w}=1$ and $\omega=1/3$ (see Eq. (2.6)). This behavior is consistent with the second equation in the VOS model (Eq. (2.7)), where setting $\bar{L}\to 0$ results in infinite curvature acceleration $k_{w}/\bar{L}\to\infty$. To ensure a finite initial velocity, $\bar{v}_{w}=1$ is imposed for the wall. However, due to resolution limitations, simulations do not explicitly capture this behavior [48]. Additionally, we note that at early times, chopping between walls can lead to an increase in $\bar{L}$, indicative of wall mergers. After reaching a maximum, chopping effectively reduces the wall size.
We observe that the domain wall rapidly expands to a superhorizon size, $\bar{L}\sim 6t$, driven by curvature acceleration resulting from the wall tension. Subsequently, the correlation length of the wall decreases due to Hubble friction, chopping effects, and friction from the bias potential. As the cosmic time approaches $t_{\rm ann}$, $\bar{L}(t)$ undergoes a significant reduction, re-entering the horizon before the domain wall annihilates at $t_{\rm ann}$. Initially, the domain walls are generated as a radiation gas, characterized by $\bar{v}_{w}=1$ and $\omega=1/3$ (see Eq. (2.6)). This behavior is consistent with the second equation in the VOS model (Eq. (2.7)), where setting $\bar{L}\to 0$ results in infinite curvature acceleration $k_{w}/\bar{L}\to\infty$. To ensure a finite initial velocity, $\bar{v}_{w}=1$ is imposed for the wall. However, due to resolution limitations, simulations do not explicitly capture this behavior [48]. Additionally, we note that at early times, chopping between walls can lead to an increase in $\bar{L}$, indicative of wall mergers. After reaching a maximum, chopping effectively reduces the wall size.
It is evident that the chopping effect should vary with time. However, the simulations in Ref. [48] assume a time-independent chopping parameter $c_{w}=0.81$, leading to the conclusion that the domain wall eventually enters the scaling regime. The chopping effect between walls can result in the efficient production of closed configurations, as confirmed by the simulations in Ref. [49]. The generation of closed domain walls could, in turn, suppress the chopping effects between walls. Once closed domain walls form at a superhorizon scale, the onset of the scaling regime for the walls could be hindered by the reduced chopping effect. In fact, the closed domain walls might never enter the scaling regime [43], contrary to previous expectations. To explore this, we adopt the following formula to describe the evolution of a spherical domain wall by neglecting the chopping effect [50]
@@
In the lower panel of Fig. 1, we plot the trajectories of the curvature radius for different initial radii, represented by red, green, and blue lines. We observe that well before the horizon crossing of the domain wall, the wall’s radius scales as $\bar{L}(t)\propto a(t)$, indicating a superhorizon size above the horizon line $\bar{L}(t)=t$. As time progresses, the pressure from the bias potential becomes significant, causing the radius trajectory to cross the horizon line at a time $t_{e}$. If $t_{e}\lesssim t_{\rm ann}$, the wall network enters the horizon before significant annihilation occurs. For the case where $t_{e}\gtrsim t_{\rm ann}$, as denoted by the blue trajectory, the annihilation occurs at the horizon crossing time $t_{e}$, rather than at $t_{\rm ann}$, leading to the phenomenon of later annihilation [43]. Thus, we confirm that for closed domain walls with negligible chopping effects, the trajectory of the curvature radius remains at a superhorizon scale and does not enter the scaling regime ($\bar{L}(t)\simeq t$) until shortly before annihilation starts.

### v3

@@ Abstract
Domain walls are topological defects produced by the spontaneous symmetry-breaking of discrete symmetry during cosmological phase transitions. Domain walls can significantly contribute to the energy density in the late-evolution stage. We propose that the density perturbations from the fluctuations in the number density of the domain walls could collapse to form primordial black holes. This mechanism becomes effective when the domain wall energy density ratio to that of the radiation reaches about 0.1 in the radiation-dominated Universe. We find that models with $Z_{2}$ symmetry are excluded for interpreting pulsar timing array observations on the nano-Hz gravitational wave background since this model’s domain wall number density fluctuations could lead to an overabundance of the primordial black holes. Moreover, the models, which generate approximately $N\sim 10$ domain walls from the spontaneous breaking of a discrete $Z_{N}$ symmetry, are also subject to stringent constraints due to the overproduction of primordial black holes.
@@ 1 Introduction
It is widely believed that new physics is required to explain various phenomena beyond the Standard Model (SM), including the neutrino mass and mixing [King:2014nza], dark matter [Bertone:2016nfn], strong CP problem [Marsh:2015xka], and matter-antimatter asymmetry [Dine:2003ax], etc. As the cosmic temperature drops with the expansion of the Universe, the spontaneous breakdown of discrete symmetries during a phase transition in new physics models can generate sheet-like topological defects, the domain walls [Zeldovich:1974uw, Kibble:1976sj]. The Universe after the phase transition is then randomly divided by them into distinct domains, which represent the clumps of the degenerate vacua. The domain walls tend to stretch in size and the neighboring domains with the same vacua merge to stabilize the network. This evolution can be described by the percolation theory [Stauffer:1978kr], whose fundamental results are based on the existence of a percolation threshold $p_{c}=0.311$ for a simple cubic lattice. The domain wall’s energy density scales as $\rho_{w}\propto t^{-1}$ in the scaling regime, which is much slower than the decay of radiation and dust, and eventually dominates the cosmological evolution. In the domain wall era, the repulsive force from domain walls would lead to an inflation phase of the Universe with the scale factor going as $a\propto t^{2}$. Such an expansion history is not compatible with the observations since it could strongly alter the Big Bang nucleosynthesis (BBN), reduce the time for the galaxy formation, and lead to unacceptably large anisotropies of the cosmic microwave background (CMB) [Planck:2015fie]. Nevertheless, the research on the nature of dark energy in the context of domain wall has been continuously conducted by recent works [Friedland:2002qs, Conversi:2004pi, Mulki:2022tfw].
@@
Discrete symmetries exist widely in new physics models. A $Z_{2}$ symmetry can emerge from the spontaneous breakdown of a gauge group in the grand unified theory (GUT) [Preskill:1992ck, Dunsky:2021tih]. The $Z_{2}$ symmetry is also commonly imposed on the two-Higgs-doublet models for the spontaneous CP violation to suppress the Higgs-mediated flavor-changing neutral currents (FCNCs) at tree level [Lee:1973iz, Preskill:1991kd, Gunion:2002zf]. The global $U(1)$ symmetry in the axion models for solving the strong CP problem is explicitly broken to the $Z_{\mathcal{N}}$ symmetry by the QCD instanton effect [Sikivie:1982qv, Preskill:1991kd]. Furthermore, to provide a dark matter candidate, a $Z_{2}$ symmetry is often imposed to make the dark matter particle stable [Chao:2017vrq, Huang:2017kzu, Grzadkowski:2018nbc, Chiang:2019oms, Chiang:2020yym, King:2023ayw, Pham:2024vso].
@@
A common solution to the domain wall problem is to introduce a bias potential $V_{\rm bias}$ to break the degenerate vacua [Dine:1993yw, Dvali:1994wv]. The domain wall network collapses as its surface energy is comparable with the bias potential, $\sigma H\simeq V_{\rm bias}$. The bias potential can be generated by the instanton effects if the global symmetry is anomalous under the framework of the theory, with such examples as the QCD and quantum gravity [Preskill:1991kd, Dine:1993yw, Dvali:1994wv, Abel:1995wk, Larsson:1996sp, Lu:2023mcz]. It is notable that, based on the successful and precise prediction in Ref. [Chiang:2020aui], Ref. [Lu:2023mcz] suggests that the recent pulsar timing array (PTA) observations, including the NANOGrav 15-year dataset (NG15) and IPTA-DR2 dataset, strongly support the mechanism of the QCD instanton-induced domain wall annihilations.
@@
Primordial black holes (PBHs) have attracted much attention since it was first suggested by Zel’dovich more than 50 years ago [Zeldovich:1967lct]. Contrary to the black holes resulting from the death and collapse of stars in their final evolution stage, PBHs are formed by cosmological processes in the early Universe. In 1971, Hawking proposed the first mechanism of PBH production [Hawking:1971ei, Carr:1974nx]: the fluctuations of a radiation-dominated Universe can produce some overdense regions. Once the density of the region $\delta=\delta\rho/\bar{\rho}$ exceeds a critical value $\delta_{c}\simeq w$, the overdense region can stop expanding and re-collapse to form PBHs when the region enters the horizon. Cosmic inflation [Ivanov:1994pa, Randall:1995dj, Garcia-Bellido:1996mdl] is commonly considered a primary source of density perturbation. The large curvature perturbations generated by the quantum fluctuations of the scalar fields are dominated by the density perturbations once they exit the horizon [Garcia-Bellido:1996mdl]. For other PBH production mechanisms, we refer to Refs. [Sasaki:2018dmp, Carr:2020gox] for a recent review.
@@
In contrast to PBH formation from the collapse of closed domain wall [Ipser:1983db, Rubin:2000dq, Ferrer:2018uiu, Liu:2019lul, Gouttenoire:2023gbn, Ge:2023rrq], we show in this work that the density perturbation due to the number density fluctuations of domain walls in the late stage could lead to the formation of PBHs. The collapse of a closed domain wall needs to satisfy the Schwarzschild criterion, i.e., the closed wall has a Schwarzschild radius near collapse [Ferrer:2018uiu]. Our scenario is based on the critical collapse of horizon-size perturbations during the horizon crossing, as outlined above. We find that if the annihilations of the domain wall are the source of the nano-Hz stochastic gravitational wave background (SGWB) observed recently by PTAs, a significant PBH relic abundance could be produced via the number density fluctuations of the walls before their annihilations.
@@ 2.1 Domain wall equation of state
The energy-momentum tensor of a plane domain wall lying in the $yz$ plane at $x=0$ (neglecting the thickness of the wall) can be estimated as [Vilenkin:1984ib, Gelmini:1988sf]
@@
where $\overline{L}=L/N$ is the average wall separation, in which there exists on average one domain wall with the surface energy-momentum tensor
@@
The average energy-momentum tensor is determined by $\overline{T}_{\mu\nu}=S_{\mu\nu}/\overline{L}$. Note that since the wall can extend along any direction, we interpret $\overline{L}$ as the correlation length of the wall. The energy density and pressure of the wall are given by the temporal and spatial components of the energy-momentum tensor
$\rho_{w}=\overline{T}_{00}=\frac{\gamma^{2}\sigma_{w}}{\overline{L}},\penalty 10000\ \penalty 10000\ {\rm and}\penalty 10000\ \penalty 10000\ p_{w}=\overline{T}_{ii}=\frac{\sigma_{w}}{3\overline{L}}(\overline{v}_{w}^{2}\gamma^{2}-2).$ (2.5)
@@ 2.2 The evolution of a domain wall
The correlation length $\bar{L}$ (or curvature radius) and the root-mean-square velocity $\bar{v}_{w}$ are two key physical quantities for describing the domain walls. We first adopt the so-called VOS model [Martins:2016book] (see appendix C for more details) to describe the evolution of $\bar{L}$ and $\bar{v}_{w}$ with cosmic expansion, which is given by
@@
where the momentum parameter $k_{w}=0.66\pm 0.04$ and the chopping parameter $c_{w}=0.81\pm 0.04$ are derived from simulations of domain wall evolution in a radiation-dominated Universe [Martins:2016ois]. The momentum parameter $k_{w}$ characterizes the curvature acceleration of the wall, while the chopping parameter $c_{w}$ represents the interaction and merger between walls. The damping length is defined as
@@
In the upper panel of Fig. 1, we illustrate the evolution of the domain wall formed at time $t_{s}$ with a symmetry-breaking scale of $v=\sigma_{w}^{1/3}=10^{5}$ GeV. The formation time of the domain wall is determined by $t_{s}=M_{\rm Pl}/(3.32g_{*}^{1/2}T_{s}^{2})$, where $g_{*}(T)\simeq 106.75$ represents the effective degrees of freedom of the plasma, $M_{\rm Pl}=1.2\times 10^{19}$ GeV is the Planck mass, and the domain wall formation temperature is assumed to be $T_{s}=v$. The initial conditions $\bar{L}(t_{s})=0.01t_{s}$ and $\bar{v}_{w}\to 0$ are adopted to solve Eqs. (2.7). We further assume a vacuum pressure $V_{\rm bias}=(10\penalty 10000\ \rm GeV)^{4}$, with the onset of the bias potential occurring at $t_{\rm on}=0.01t_{\rm ann}$.
The red curves illustrate the evolution of $\bar{L}(t)$, while the blue curves denote that of $\bar{v}_{w}$. To estimate the impact of the chopping effect, the chopping parameter $c_{w}$ is set to 0.81 and 0.61 for the solid and dashed curves, respectively. We summarize the key observational results from Fig. 1 as follows:
  • Initially, domain walls are produced as a radiation-like fluid, exhibiting $\bar{v}_{w}=1$ and $\omega=1/3$ (see Eq. (2.6)). This dynamical feature is consistent with the velocity evolution equation in the VOS model (Eq. (2.7)): when $\bar{L}\to 0$, the curvature acceleration $k_{w}/\bar{L}\to\infty$, which would otherwise lead to an unphysical infinite velocity. To enforce a finite initial velocity, we impose $\bar{v}_{w}=1$ for the domain walls so that $d\bar{v}_{w}/dt=0$. However, due to limitations in numerical resolution, simulations in Ref. [Martins:2016ois] fail to explicitly capture this initial relativistic behavior.
  • Driven by curvature acceleration induced by wall tension, the domain walls rapidly expand to a superhorizon scale with $\bar{L}\sim 6t$. To clarify this evolution, we fix $\bar{v}_{w}=1$ in the first equation of the VOS model (Eq. (2.7)) before $\bar{L}/t$ attains its maximum value. Prior to the activation of the bias potential, the damping length satisfies $1/\ell_{d}=3H$; combined with the radiation-dominated Universe condition $H=1/(2t)$, the characteristic scale of the domain walls initially grows as $\bar{L}\propto t^{2}$. As $\bar{L}$ increases, the curvature acceleration gradually weakens, while the relativistic speed of the domain walls can be sustained until the curvature acceleration is overtaken by Hubble damping. This transition occurs when the characteristic scale reaches
    $\bar{L}=\frac{6k_{w}}{\bar{v}_{w}}t\sim 6t\penalty 10000\ ,$ (2.10)
    which is determined by the balance condition $k_{w}/\bar{L}=\bar{v}_{w}/\ell_{d}$. Notably, owing to the wall’s curvature acceleration, domain walls can retain extreme relativistic speeds ($\bar{v}_{w}\simeq 1$) even after their characteristic scale enters the superhorizon regime (i.e., $\bar{L}\gtrsim t$).
  • Once the domain walls reach the superhorizon scale of $\bar{L}\sim 6t$, the condition $k_{w}/\bar{L}\lesssim\bar{v}_{w}/\ell_{d}$ is satisfied, meaning that the local deceleration due to friction becomes more dominant than curvature acceleration. Subsequently, the correlation length of the domain walls decreases to the horizon scale under the combined influence of Hubble friction, chopping effects, and friction from the bias potential. As cosmic time approaches $t_{\rm ann}$, $\bar{L}(t)$ experiences a substantial decrease, re-entering the horizon prior to the domain walls’ annihilation at $t_{\rm ann}$.
  • Furthermore, we note that at the early stages, chopping between domain walls can cause an increase in $\bar{L}$, which is a signature of domain wall mergers. Following the attainment of a maximum value, chopping effectively reduces the domain wall scale.
It is important to emphasize that the curvature acceleration of domain walls is a global mechanical effect induced by wall tension—analogous to the stretching of an elastic membrane—whereby each segment of the wall undergoes synchronous stretching under the tension, independent of local information propagation. Consequently, the curvature radius, or equivalently the characteristic scale of the domain wall, can extend to a superhorizon regime ($\bar{L}\sim 6t$) without violating causality constraints. In contrast, the propagation speed of damping effects, e.g., the Hubble friction, cannot exceed the speed of light and thus is subject to causality limitations. As a result, the characteristic length associated with damping effects must not be larger than the Hubble scale. For the Hubble friction term in Eq. (2.7), the damping length is $\ell_{d}=1/(3H)=2t/3$, which is explicitly subhorizon.
In the initial phase, the curvature acceleration dominates the expansion of domain walls, driving them to superhorizon scales while sustaining extremely relativistic speeds ($\bar{v}_{w}\simeq 1$). As the characteristic scale of the domain walls expands to $\bar{L}\sim 6t$, the efficiency of curvature acceleration diminishes significantly. Both friction and chopping are local effects; once curvature acceleration weakens, these local processes take over the evolution of the superhorizon-sized domain walls, ultimately leading to their re-entry into the horizon.
The primary role of the chopping effect is not to maintain the scaling regime, but to regulate the transition into it by controlling when superhorizon-sized walls re-enter the horizon. We therefore argue that its key significance lies in setting the timing of this horizon re-entry. As shown below, this timing directly determines the amplitude of the resultant density perturbation, and consequently the abundance of PBHs. In contrast, consistent with simulations [Martins:2016ois, Dankovsky:2024zvs], chopping is not an essential mechanism for sustaining the scaling regime itself. While the chopping effect is expected to be time-dependent, the analysis in Ref. [Martins:2016ois] employs a constant chopping parameter for simplicity. The chopping effect between walls can result in the efficient production of closed configurations, as confirmed by the simulations [Lazanu:2015fua]. The generation of closed domain walls could, in turn, suppress the chopping effects between walls. Once closed domain walls form at a superhorizon scale, the onset of the scaling regime for the walls could be hindered by the reduced chopping effect. It has been suggested that superhorizon-sized closed domain walls might never attain the scaling regime prior to the onset of appreciable pressure from the bias potential [Gouttenoire:2023gbn]. To explore this, we adopt the following formula to describe the evolution of a spherical domain wall by neglecting the chopping effect [Deng:2020mds]
@@
In the lower panel of Fig. 1, we plot the trajectories of the curvature radius for different initial radii, represented by red, green, and blue lines. We observe that well before the horizon crossing of the domain wall, the wall’s radius scales as $\bar{L}(t)\propto a(t)$, indicating a superhorizon size above the horizon line $\bar{L}(t)=t$. As time progresses, the pressure from the bias potential becomes significant, causing the radius trajectory to cross the horizon line at a time $t_{e}$. If $t_{e}\lesssim t_{\rm ann}$, the wall network enters the horizon before significant annihilation occurs. For the case where $t_{e}\gtrsim t_{\rm ann}$, as denoted by the blue trajectory, the annihilation occurs at the horizon crossing time $t_{e}$, rather than at $t_{\rm ann}$, leading to the phenomenon of later annihilation [Gouttenoire:2023gbn]. Thus, we confirm that for closed domain walls with negligible chopping effects, the trajectory of the curvature radius remains at a superhorizon scale and does not enter the scaling regime ($\bar{L}(t)\simeq t$) until shortly before annihilation starts.